OpenStudy (anonymous):

One leg of a triangle has a measure of 24 units. If the hypotenuse measures 25 units, what is the length of the other leg?

OpenStudy (pigrenex):

\[25^{2} = 24^{2} + x ^{2}\] \[25^{2} - 24^{2} = x ^{2}\] \[625 - 576 = x^2\] \[49 = x^2\] \[x = \sqrt{49}\] x = 7

OpenStudy (anonymous):

thank you again!LOL

OpenStudy (pigrenex):

no problem :D